OpenStudy (linna):

find volume of regular pyramid if is given: area of the base 12 cm^2 , and area of 4 sides is 24 cm^2 . find Volume

OpenStudy (linna):

Pyramid looks like http://math.about.com/od/formulas/ss/surfaceareavol_5.htm

OpenStudy (linna):

multiple choices are 10 cm^3 12 16 20

OpenStudy (savvy):

12cm3

OpenStudy (linna):

I want to know how to do it, I'm trying to solve this but I can't ...

OpenStudy (linna):

could it be something like: \[ \large \begin{array}{l}A = 2bs + {b^2}\\24 + 14 = 2bs + 12\\36 - 12 = 2bs\\\frac{{24}}{2} = bs\\12 = \sqrt {12} \cdot s\\s = \frac{{12}}{{\sqrt {12} }}\end{array} \]

jhonyy9 (jhonyy9):

,,Linna" can you tel me what form sides has a pyramid ?

jhonyy9 (jhonyy9):

24cm^2 is the total lateral area ?

OpenStudy (linna):

is like this: (is not drawn in the given, but this is I'm sure... ) http://math.about.com/od/formulas/ss/surfaceareavol_5.htm

OpenStudy (savvy):

\[4\times(1/2)sb = 24\]b^2 = 12 from the above eqn's we get s=b h = sqrt.[s^2 - (b/2)^2] since s=b; h = 3 hence volume = (1/2)X3X12 =12

OpenStudy (linna):

yes lateral that's the word i couldn't find ... is 24 cm^2 total lateral so one it means 24/4 right? ..

jhonyy9 (jhonyy9):

yes sure

jhonyy9 (jhonyy9):

so there on this site what you have attached are all dates what you need for you can solveing this exercise right

jhonyy9 (jhonyy9):

what the base of a pyramid ?

OpenStudy (linna):

square... 

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

than area of a square = ?

OpenStudy (linna):

 \[\ a\times a=a^2 \]